A polynomials is defined by P(x) = mx³ + mx² + mx + n where m and n are non-zero real constants. Given that x = -n is a root of P(x). Determine the range of the possible values of m.

1
Expert's answer
2021-11-08T16:57:04-0500
"P(-n)=-mn^3+mn^2-mn+n=0"

"n\not=0"


"-mn^2+mn-m+1=0"

"m=\dfrac{1}{n^2-n+1}"

Let "m=f(n)=\dfrac{1}{n^2-n+1}, mn\not=0"


"n^2-n+1=n^2-n+\dfrac{1}{4}+\dfrac{3}{4}"

"=(n-\dfrac{1}{2})^2+\dfrac{3}{4}>0,n\in \R"

"f'(n)=-\dfrac{2n-1}{(n^2-n+1)^2}"

Find the critical number(s)


"f'(n)=0=>-\dfrac{2n-1}{(n^2-n+1)^2}=0=>n=\dfrac{1}{2}"

If "n\in(-\infin, 0)\cup(0, \dfrac{1}{2})," then "f'(n)>0, f(n)" increases.

If "n\in(\dfrac{1}{2},\infin)," then "f'(n)<0, f(n)" decreases.

The function "f(n)" has a local maximum at "n=\dfrac{1}{2}."


"f(\dfrac{1}{2})=\dfrac{1}{(\dfrac{1}{2})^2-\dfrac{1}{2}+1}=\dfrac{4}{3}"

"\lim\limits_{n\to0^-}f(n)=\lim\limits_{n\to0^+}f(n)=1"

"\lim\limits_{n\to-\infin}f(n)=\lim\limits_{n\to\infin}f(n)=0"

Then "m\in(0, \dfrac{4}{3}]"